# Automating Android Build Repair: Bridging the Reasoning-Execution Gap in LLM Agents with Domain-Specific Tools

### v2

@@ F.2.2 Gemini-CLI with Shell (Failed Trajectory)
Failed Trajectory: Gemini-CLI with Shell Step 1: Get Detailed Error Information Thought: The build failed with a Kapt error. I need more information. I’$l$ use the shell to run the build with the --stacktrace flag as suggested. Tool Call: run_shell() Step 2: Identify the Root Cause and Apply a Plausible Fix Thought: The stack trace clearly shows a data binding error in fragment_about.xml caused by a type mismatch on a ‘TextView‘. The line has a ‘default=0‘ integer value for a string attribute. This is the obvious error. I will remove the default value. Tool Call: replace(..., old_string=’android:text="@{context.version, default=0}"’, new_string=’android:text="@{context.version}"’) Step 3: Abandon the Correct Path Thought: I’$l$ rebuild to verify my fix. Tool Call: run_shell(shell_command=) Result: BUILD FAILED. The same data binding error persists. My fix didn’t work. I must have been wrong. The *real* problem must be the Kotlin version warning I saw in the first log. I will ignore the data binding error and focus on fixing the build environment. Step 4: Descend into "Version Hell" Thought: The project uses Kotlin ‘2.0.20‘, which Kapt doesn’t support. This must be the root cause. I will downgrade everything. I will downgrade the Kotlin version, the Compose plugin version, and the Compose BOM version to older, more stable releases. Tool Call: replace(file_path=’.../build.gradle.kts’, old_string=’val kotlinVersion = "2.0.20"’, new_string=’val kotlinVersion = "1.9.22"’) Tool Call: replace(file_path=’.../app/build.gradle.kts’, old_string=’...compose") version "2.0.20"’, new_string=’...compose") version "1.5.10"’) Step 5: Create New Errors Thought: I have downgraded the versions. I will try building again. Tool Call: run_shell(shell_command=) Result: BUILD FAILED. A new error appears: Plugin [id: org.jetbrains.kotlin.plugin.compose’, version: ’1.5.10’] was not found. My attempt to fix the versions has made things worse. Step 6: Failure Thought: I am stuck in a loop. Every change I make to the versioning either fails or creates a new problem. My attempts to fix the build environment have been fruitless. I have exhausted my strategies. Result: The agent terminates after reaching its maximum number of attempts. It never solved the problem because it abandoned the correct debugging path and got lost chasing a red herring.
F.3 Case 3: Unresolved Dependency and Flawed Reasoning
@@
F.3.1 GradleFixer (Successful Trajectory)
@@
Successful Trajectory: GradleFixer Step 1: Diagnose the Missing Dependency Thought: The build is failing with many "Unresolved reference" errors pointing to ‘MediaMetadata.kt‘. The missing references like ‘ImageMetadataReader‘ and ‘ExifIFD0Directory‘ suggest a problem with an image metadata extraction library. I will investigate the project’s dependencies to resolve this. Tool Call: read_file(path=’Gallery/gradle/libs.versions.toml’) Step 2: Identify Dependency Management Strategy Thought: The project uses a TOML version catalog. I’ve confirmed that the ‘metadata-extractor‘ library is not defined. I need to add it to ‘libs.versions.toml‘ and then include it in the app’s build script. Tool Call: read_file(path=’Gallery/app/build.gradle.kts’) Step 3: Add Dependency to Version Catalog Thought: I will now add the ‘metadata-extractor‘ library to the ‘libs.versions.toml‘ file. First, I’$l$ add the version to the ‘[versions]‘ section, then the library definition to the ‘[libraries]‘ section to maintain consistency. Tool Call: replace(file_path=’Gallery/gradle/libs.versions.toml’, old_string=’...’, new_string=’...’) Tool Call: replace(file_path=’Gallery/gradle/libs.versions.toml’, old_string=’...’, new_string=’...’) Step 4: Add Dependency to App Module Thought: Now that the library is defined in the version catalog, I will add it as an ‘implementation‘ dependency in the ‘app/build.gradle.kts‘ file. Tool Call: replace(file_path=’Gallery/app/build.gradle.kts’, old_string=’...’, new_string=’...’) Step 5: Verify the Fix Thought: The dependency has been added to the project configuration. I will now run the build to confirm that the issue is resolved. Tool Call: TOOL_A() Result: BUILD SUCCESSFUL. The missing dependency was the root cause. The task is complete.
F.3.2 Gemini-CLI with Shell (Failed Trajectory)
@@
Failed Trajectory: Gemini-CLI with Shell Step 1: Initial Diagnosis and Correct Fix Thought: The errors in ‘MediaMetadata.kt‘ indicate a missing dependency. The code uses classes from the ‘com.drew…‘ package, so I need to add the ‘metadata-extractor‘ library. I will add it to ‘libs.versions.toml‘ and ‘app/build.gradle.kts‘. Tool Call: (Agent successfully adds the dependency to both files.) Step 2: Flawed Reasoning and Reversal Thought: I’ve added the ‘metadata-extractor‘ dependency to fix the issue. I’$l$ now remove the artifacts made to the build files. Tool Call: (Agent reverts its own correct changes from both ‘app/build.gradle.kts‘ and ‘gradle/libs.versions.toml‘.) Step 3: Unnecessary Cache Cleaning Thought: I reverted the changes. Now, I will try to build again. (Build fails with the original error). The build failed again. This often indicates a corrupted cache. I’$l$ clean the build cache and try building again. Tool Call: run_shell(shell_command=) Result: Command fails as ‘cleanBuildCache‘ is not a valid task. The agent then tries ‘clean‘. Step 4: Rediscovery and Repeated Failures Thought: It appears I mistakenly reverted the fix that should have resolved the build error. My apologies. The "Unresolved reference" errors indicate that the ‘metadata-extractor‘ dependency is indeed required. I will now re-apply the fix. Tool Call: (Agent now enters a long loop of failures, repeatedly attempting to modify the Gradle files but failing due to incorrect syntax for referencing version catalog aliases with hyphens, e.g., using ‘libs.metadata.extractor‘ instead of the correct ‘libs.metadataExtractor‘.) Step 5: Failure Thought: My apologies for the repeated mistakes. It seems I’m stuck in a loop. I will try to fix the issue by first re-reading the ‘gradle/libs.versions.toml‘ file to confirm the correct dependency alias. Result: The agent continues to struggle with the correct syntax and file paths, unable to break out of its error loop. It eventually terminates after reaching its maximum number of attempts, having failed to re-apply the fix it had already discovered once.
Appendix G Curated Projects in AndroidBuildBench

### v3

@@ E.2.2 Gemini-CLI with Shell (Failed Trajectory)
Failed Trajectory: Gemini-CLI with Shell Step 1: Get Detailed Error Information Thought: The build failed with a Kapt error. I need more information. I’$l$ use the shell to run the build with the --stacktrace flag as suggested. Tool Call: run_shell(shell_command=./gradlew assembleDebug --stacktrace’) Step 2: Identify the Root Cause and Apply a Plausible Fix Thought: The stack trace clearly shows a data binding error in fragment_about.xml caused by a type mismatch on a ‘TextView‘. The line has a ‘default=0‘ integer value for a string attribute. This is the obvious error. I will remove the default value. Tool Call: replace(..., old_string=android:text="@{context.version, default=0}"’, new_string=android:text="@{context.version}"’) Step 3: Abandon the Correct Path Thought: I’$l$ rebuild to verify my fix. Tool Call: run_shell(shell_command=./gradlew assembleDebug --parallel’) Result: BUILD FAILED. The same data binding error persists. My fix didn’t work. I must have been wrong. The *real* problem must be the Kotlin version warning I saw in the first log. I will ignore the data binding error and focus on fixing the build environment. Step 4: Descend into "Version Hell" Thought: The project uses Kotlin ‘2.0.20‘, which Kapt doesn’t support. This must be the root cause. I will downgrade everything. I will downgrade the Kotlin version, the Compose plugin version, and the Compose BOM version to older, more stable releases. Tool Call: replace(file_path=.../build.gradle.kts’, old_string=val kotlinVersion = "2.0.20"’, new_string=val kotlinVersion = "1.9.22"’) Tool Call: replace(file_path=.../app/build.gradle.kts’, old_string=...compose") version "2.0.20"’, new_string=...compose") version "1.5.10"’) Step 5: Create New Errors Thought: I have downgraded the versions. I will try building again. Tool Call: run_shell(shell_command=./gradlew assembleDebug --parallel’) Result: BUILD FAILED. A new error appears: Plugin [id: org.jetbrains.kotlin.plugin.compose’, version: ’1.5.10’] was not found. My attempt to fix the versions has made things worse. Step 6: Failure Thought: I am stuck in a loop. Every change I make to the versioning either fails or creates a new problem. My attempts to fix the build environment have been fruitless. I have exhausted my strategies. Result: The agent terminates after reaching its maximum number of attempts. It never solved the problem because it abandoned the correct debugging path and got lost chasing a red herring.
E.3 Case 3: Unresolved Dependency and Flawed Reasoning
@@
E.3.1 GradleFixer (Successful Trajectory)
@@
Successful Trajectory: GradleFixer Step 1: Diagnose the Missing Dependency Thought: The build is failing with many "Unresolved reference" errors pointing to ‘MediaMetadata.kt‘. The missing references like ‘ImageMetadataReader‘ and ‘ExifIFD0Directory‘ suggest a problem with an image metadata extraction library. I will investigate the project’s dependencies to resolve this. Tool Call: read_file(path=Gallery/gradle/libs.versions.toml’) Step 2: Identify Dependency Management Strategy Thought: The project uses a TOML version catalog. I’ve confirmed that the ‘metadata-extractor‘ library is not defined. I need to add it to ‘libs.versions.toml‘ and then include it in the app’s build script. Tool Call: read_file(path=Gallery/app/build.gradle.kts’) Step 3: Add Dependency to Version Catalog Thought: I will now add the ‘metadata-extractor‘ library to the ‘libs.versions.toml‘ file. First, I’$l$ add the version to the ‘[versions]‘ section, then the library definition to the ‘[libraries]‘ section to maintain consistency. Tool Call: replace(file_path=Gallery/gradle/libs.versions.toml’, old_string=...’, new_string=...’) Tool Call: replace(file_path=Gallery/gradle/libs.versions.toml’, old_string=...’, new_string=...’) Step 4: Add Dependency to App Module Thought: Now that the library is defined in the version catalog, I will add it as an ‘implementation‘ dependency in the ‘app/build.gradle.kts‘ file. Tool Call: replace(file_path=Gallery/app/build.gradle.kts’, old_string=...’, new_string=...’) Step 5: Verify the Fix Thought: The dependency has been added to the project configuration. I will now run the build to confirm that the issue is resolved. Tool Call: run_build() Result: BUILD SUCCESSFUL. The missing dependency was the root cause. The task is complete.
E.3.2 Gemini-CLI with Shell (Failed Trajectory)
@@
Failed Trajectory: Gemini-CLI with Shell Step 1: Initial Diagnosis and Correct Fix Thought: The errors in ‘MediaMetadata.kt‘ indicate a missing dependency. The code uses classes from the ‘com.drew…‘ package, so I need to add the ‘metadata-extractor‘ library. I will add it to ‘libs.versions.toml‘ and ‘app/build.gradle.kts‘. Tool Call: (Agent successfully adds the dependency to both files.) Step 2: Flawed Reasoning and Reversal Thought: I’ve added the ‘metadata-extractor‘ dependency to fix the issue. I’$l$ now remove the artifacts made to the build files. Tool Call: (Agent reverts its own correct changes from both ‘app/build.gradle.kts‘ and ‘gradle/libs.versions.toml‘.) Step 3: Unnecessary Cache Cleaning Thought: I reverted the changes. Now, I will try to build again. (Build fails with the original error). The build failed again. This often indicates a corrupted cache. I’$l$ clean the build cache and try building again. Tool Call: run_shell(shell_command=./gradlew cleanBuildCache’) Result: Command fails as ‘cleanBuildCache‘ is not a valid task. The agent then tries ‘clean‘. Step 4: Rediscovery and Repeated Failures Thought: It appears I mistakenly reverted the fix that should have resolved the build error. My apologies. The "Unresolved reference" errors indicate that the ‘metadata-extractor‘ dependency is indeed required. I will now re-apply the fix. Tool Call: (Agent now enters a long loop of failures, repeatedly attempting to modify the Gradle files but failing due to incorrect syntax for referencing version catalog aliases with hyphens, e.g., using ‘libs.metadata.extractor‘ instead of the correct ‘libs.metadataExtractor‘.) Step 5: Failure Thought: My apologies for the repeated mistakes. It seems I’m stuck in a loop. I will try to fix the issue by first re-reading the ‘gradle/libs.versions.toml‘ file to confirm the correct dependency alias. Result: The agent continues to struggle with the correct syntax and file paths, unable to break out of its error loop. It eventually terminates after reaching its maximum number of attempts, having failed to re-apply the fix it had already discovered once.
Appendix F Curated Projects in AndroidBuildBench